# Causal Schrödinger Bridges: Constrained Optimal Transport on Structural Manifolds

### v2

@@ Abstract
Generative modeling typically seeks the path of least action via deterministic flows (ODE). While effective for in-distribution tasks, we argue that these deterministic paths become brittle under causal interventions, which often require transporting probability mass across low-density regions ("off-manifold") where the vector field is ill-defined. This leads to numerical instability and spurious correlations. In this work, we introduce the Causal Schrödinger Bridge (CSB), a framework that reformulates counterfactual inference as Entropic Optimal Transport. Unlike deterministic approaches that require strict invertibility, CSB leverages diffusion processes (SDEs) to robustly "tunnel" through support mismatches while strictly enforcing structural admissibility constraints. We prove the Structural Decomposition Theorem, showing that the global high-dimensional bridge factorizes exactly into local, robust transitions. Remarkably, our framework achieves linear scalability, solving a 1000-dimensional causal transport problem in just 14 seconds on a single consumer-grade GPU(3090). Empirical validation on Morpho-MNIST and 1000-D stress tests demonstrates that CSB significantly outperforms deterministic baselines in structural consistency and distribution coverage, particularly in regimes of strong, out-of-distribution treatments. Our code is available at: https://github.com/cochran1/causal-schrodinger-bridge.
@@ 1 Introduction
A strong causal intervention (e.g., $do(T=t^{\prime})$) often necessitates transporting a sample from the factual distribution $P(X|t)$ to a counterfactual region $P(X|t^{\prime})$ that may be topologically disjoint or separated by low-density "voids" (off-manifold regions). In these regimes, deterministic flows often become numerically unstable or find energy-minimizing "shortcuts" that violate causal laws. We term this phenomenon Anticipatory Control: future effects implicitly steering past causes to minimize global transport cost. While exact inversion can theoretically mitigate this [9], numerical instability in low-density regions often re-introduces these artifacts in practice.
@@
    Geometric Framework: We formalize counterfactual generation as a constrained entropy-regularized optimal transport problem, identifying the geometric conflict between manifold adherence (Least Action) and structural adherence (Arrow of Time).
@@
    Theory & Extreme Scalability: We prove the Structural Decomposition Theorem (Theorem 1), showing that the intractable high-dimensional bridge problem factorizes exactly into a sequence of local, robust transitions. Crucially, this factorization bypasses the curse of dimensionality, enabling linear scalability $O(d)$ for causal transport.
@@
    High-Dimensional Empirical Benchmarks: We validate CSB on challenging high-dimensional tasks:
    • The 1000-D Barrier: In a 1000-dimensional "causal surgery" stress test, CSB completes training in just 14 seconds on a single consumer GPU, achieving full support coverage where deterministic baselines suffer from 18% mode collapse.
    • Mechanism Disentanglement: On Morpho-MNIST (784-D), CSB disentangles causal factors (thickness) from style without supervision, achieving a 4.5x improvement in structural consistency over state-of-the-art baselines.
@@ 2.1 Schrödinger Bridges (SB)
The standard SB problem seeks a path measure $\mathbb{P}$ on path space $\Omega=C([0,1],\mathbb{R}^{d})$ that couples $\mu_{0},\mu_{1}$ while minimizing the Kullback-Leibler divergence w.r.t. a reference measure $\mathbb{Q}$ [2]:
$\min_{\mathbb{P}\in\mathcal{P}(\Omega)}\mathrm{KL}(\mathbb{P}||\mathbb{Q})\quad\text{s.t.}\quad P_{0}=\mu_{0},P_{1}=\mu_{1}.$ (1)
@@ 3.1 The Factorized Reference Process
The reference measure $\mathbb{Q}$ is the law of the solution to the SDE system:
@@ 3.3 The Causal Schrödinger Bridge Problem
We formulate the CSB as a constrained optimization problem. We seek the measure that is closest to the reference $\mathbb{Q}$ while satisfying boundary conditions AND the causal admissibility constraint.
@@
$\mathbb{P}^{*}=\arg\min_{\mathbb{P}\in\mathcal{P}_{\mathcal{G}}}\mathrm{KL}(\mathbb{P}||\mathbb{Q})\quad\text{s.t.}\quad P_{0}=\mu_{0},P_{1}=\mu_{1}.$ (3)
@@ 3.4 The Structural Decomposition Theorem
Remark 1 (Geometric Interpretation: Entropic Tunneling vs. Deterministic Rigidity).
It is crucial to clarify that the diffusion term $g(t)\mathrm{d}\mathbf{W}_{t}$ in our framework is not an ad-hoc heuristic for robustness, but the necessary consequence of Entropic Regularization in the Schrödinger Bridge formulation. Geometrically, standard deterministic transport (ODE) is forced to follow geodesics on the statistical manifold, often failing to traverse low-density regions ("voids") caused by support mismatch. The entropic term "relaxes" this rigidity, allowing the optimal transport plan to "tunnel" through these voids via Brownian motion while maintaining the distributional envelope. Crucially, Theorem 1 guarantees that this stochastic exploration strictly adheres to the causal graph, effectively decomposing the high-dimensional tunneling problem into local, structurally valid transitions.
@@ 4.3 Geometric Stress Test: Tunneling through the High-Dimensional Void
While the theoretical Causal Schrödinger Bridge implies exact entropic minimization, in high-dimensional tasks ($D\geq 50$), we employ a Robust Conditional Flow Matching (CFM) objective for training stability. Specifically, we inject Gaussian noise during training to ensure the vector field is robust to off-manifold perturbations.
@@ 4.4 The 1000-D Barrier: Causal Precision at Scale
Quantitative Stress Test ($D=1000$): To demonstrate the linear scalability promised by Theorem 1, we simulate a 1000-dimensional causal surgery. Standard entropic OT methods typically fail to converge at this scale.
@@
Total Training Time $\downarrow$ 13.87 s 13.65 s Linear Scaling
@@
Addressing the Metric Paradox: As shown in Table 2, while ODE appears more "precise" in leakage (0.03 vs 0.19), this is a classic manifestation of Mode Collapse. The ODE baseline regresses to a rigid, zero-entropy line, failing to cover 18% of the target support. CSB chooses distributional authenticity over the "safe" mean-path regression. We argue that a structurally valid probabilistic envelope is far more valuable than a "precisely wrong" deterministic trajectory[cite: 143].
@@ 5 Discussion: From Digits to Discovery
Scaling the Invisible Wall: Beyond 1000 Dimensions. A pivotal contribution of this work is the demonstration that structural intelligence can bypass the curse of dimensionality. As evidenced by our 1000-D causal surgical test (Section 4.4), CSB maintains linear scalability, completing a complex high-dimensional transport task in just 14 seconds on a single consumer-grade GPU(3090). This efficiency is the direct empirical manifestation of our Structural Decomposition Theorem , proving that when a generative model respects the causal backbone, the computational complexity is no longer an exponential barrier but a manageable linear overhead.
Computational Democratization in Causal AI. This efficiency marks a shift towards the democratization of high-stakes causal inference. By reducing the computational cost of 1000-node system modeling from high-performance clusters to individual researchers’ desktops, CSB enables rapid, iterative hypothesis testing in fields where data is high-dimensional but compute may be constrained.
Implications for AI for Science. While we validated CSB on visual proxies like Morpho-MNIST, the framework is structurally isomorphic to the most pressing challenges in Single-cell Perturbation Prediction (e.g., Perturb-seq [5]). In genomics, gene knockouts ($T$) causally influence high-dimensional expression states ($\mathbf{X}$) via complex regulatory networks. Standard OT often traverses biologically impossible "hybrid" states because it ignores the structural manifold. Our 1000-D results provide a rigorous foundation for scaling CSB to whole-genome regulatory networks, moving beyond "black-box" generation to mechanistically valid discovery tools.
Limitations and Future Work. Currently, CSB assumes a known DAG[cite: 303]. Future research will focus on integrating Differentiable Causal Discovery to simultaneously learn the graph topology and the transport map, potentially enabling the autonomous discovery of mechanisms from raw observational trajectories.
@@ Appendix A Appendix: Geometric Interpretations
Figure 7: The Causal Light Cone vs. Anticipatory Control. Visualizing transport paths in space-time $(X,t)$. Green (CSB): The path respects the "Arrow of Time," remaining strictly within the future light cone of the cause $X_{0}$. Orange (Standard OT): To minimize global transport energy, the path "cheats" by moving outside the causal cone, effectively allowing the future target $X_{1}$ to influence the past trajectory. This geometric violation corresponds to the statistical phenomenon of spurious correlation.
@@ B.1 Notation and Measure Theoretic Setup
Let $\Omega=C([0,1],\mathbb{R}^{d})$ be the path space equipped with the canonical filtration $\mathcal{F}_{t}=\sigma(\mathbf{X}_{s}:s\leq t)$. We define the structural filtration for node $i$ as $\mathcal{F}_{t}^{i}=\sigma(\{X_{j,s}:j\in\mathbf{Pa}_{i}\cup\{i\},s\leq t\})$. The reference measure $\mathbb{Q}$ is induced by the SDE system where $f_{i}$ depends only on parents:
@@ B.2 Proof of Theorem 1 (Structural Decomposition)
$\min_{\mathbb{P}\in\mathcal{P}_{\mathcal{G}}}\mathrm{KL}(\mathbb{P}||\mathbb{Q})\quad\text{s.t.}\quad P_{0}=\mu_{0},P_{1}=\mu_{1}.$ (11)

### v3

@@ Abstract
Generative modeling typically seeks the path of least action via deterministic flows (ODE). While effective for in-distribution tasks, we argue that these deterministic paths become brittle under causal interventions, which often require transporting probability mass across low-density regions (“off-manifold”) where the vector field is ill-defined. This leads to numerical instability and spurious correlations. In this work, we introduce the Causal Schrödinger Bridge (CSB), a framework that reformulates counterfactual inference as Entropic Optimal Transport. Unlike deterministic approaches that require strict invertibility or rely on low-rank approximations, CSB leverages diffusion processes (SDEs) to robustly “tunnel” through support mismatches while strictly enforcing structural admissibility constraints. We prove the Structural Decomposition Theorem, showing that the global high-dimensional bridge factorizes exactly into local, robust transitions.
Crucially, we demonstrate that CSB breaks the Curse of Dimensionality in regimes of high intrinsic dimension. We empirically validate this on a full-rank causal system ($d=10^{5}$, intrinsic rank $10^{5}$), completing the transport in 26.48 seconds on a single GPU(3090). This stands in stark contrast to structure-agnostic $O(d^{3})$ baselines, which are estimated to require over 6 years for dense computations of this scale regardless of the data’s intrinsic rank. Empirical validation on Morpho-MNIST and $10^{5}$-D extremal stress tests demonstrates that CSB significantly outperforms deterministic baselines in structural consistency and distribution coverage, capturing the underlying manifold with high fidelity ($MSE\approx 0.04$). Our code is available at: https://github.com/cochran1/causal-schrodinger-bridge.
@@ 1 Introduction
A strong causal intervention (e.g., $do(T=t^{\prime})$) often necessitates transporting a sample from the factual distribution $P(\mathbf{X}|t)$ to a counterfactual region $P(\mathbf{X}|t^{\prime})$ that may be topologically disjoint or separated by low-density "voids" (off-manifold regions). In these regimes, deterministic flows often become numerically unstable or find energy-minimizing "shortcuts" that violate causal laws. We term this phenomenon Anticipatory Control: future effects implicitly steering past causes to minimize global transport cost. While exact inversion can theoretically mitigate this [9], numerical instability in low-density regions often re-introduces these artifacts in practice.
@@
    Geometric Framework & Hybrid Strategy: We identify the fundamental tension between Identity Preservation (requiring determinism) and Robust Transport (requiring stochasticity) in causal inference. We propose a Hybrid Inference Strategy that resolves this paradox: using deterministic ODEs for precise abduction and entropic SDEs for robust counterfactual tunneling.
@@
    Extremal Scalability in Full-Rank Regimes: We address the Curse of Dimensionality via structural decomposition. We first calibrate the cubic cost of standard dense solvers ($O(d^{3})$), showing they remain prohibitive ($\approx 6.37$ years) for high-dimensional inputs regardless of rank. We then confront CSB with a Full-Rank Stress Test ($d=10^{5}$), where every dimension is causally active ($X_{t}\leftarrow X_{t-1}$). CSB completes this task in 26.48 seconds, achieving a speedup of order $10^{7}$ over the extrapolated baseline and proving linear scalability $O(d)$ even when the system allows no low-dimensional embedding.
@@
    High-Dimensional Benchmarks: On Morpho-MNIST (784-D), our hybrid CSB disentangles causal factors from style without supervision, achieving a 3.2x improvement in structural consistency (SSIM) and a 72% reduction in transport cost compared to state-of-the-art deterministic baselines.
@@ 2.1 Schrödinger Bridges (SB)
The standard SB problem seeks a path measure $\mathbb{P}$ on path space $\Omega=C([0,1],\mathbb{R}^{d})$ that couples the marginal distributions $\mu_{0},\mu_{1}$ while minimizing the Kullback-Leibler divergence w.r.t. a reference measure $\mathbb{Q}$ [2]:
$\min_{\mathbb{P}\in\mathcal{P}(\Omega)}\mathrm{KL}(\mathbb{P}||\mathbb{Q})\quad\text{s.t.}\quad\mathbb{P}_{0}=\mu_{0},\mathbb{P}_{1}=\mu_{1}.$ (1)
@@ 3.1 The Factorized Reference Process
The reference measure $\mathbb{Q}\in\mathcal{P}(\Omega)$ is the law of the solution to the SDE system:
@@ 3.3 The Causal Schrödinger Bridge Problem
We formulate the CSB as a constrained optimization problem. We seek the path measure that is closest to the reference $\mathbb{Q}$ while satisfying boundary conditions AND the causal admissibility constraint.
@@
$\mathbb{P}^{*}=\arg\min_{\mathbb{P}\in\mathcal{P}_{\mathcal{G}}}\mathrm{KL}(\mathbb{P}||\mathbb{Q})\quad\text{s.t.}\quad\mathbb{P}_{0}=\mu_{0},\mathbb{P}_{1}=\mu_{1}.$ (3)
@@ 3.4 The Structural Decomposition Theorem
Remark 1 (Geometric Interpretation: Entropic Tunneling vs. Scalability).
It is crucial to clarify that the diffusion term $g(t)\mathrm{d}\mathbf{W}_{t}$ in our framework is not an ad-hoc heuristic for robustness, but the necessary consequence of Entropic Regularization in the Schrödinger Bridge formulation. Geometrically, standard deterministic transport (ODE) is forced to follow geodesics on the statistical manifold, often failing to traverse low-density regions (“voids”) caused by support mismatch. The entropic term “relaxes” this rigidity, allowing the transport plan to “tunnel” through these voids via Brownian motion. Crucially, this geometric relaxation is the key to our linear scalability $O(d)$. While deterministic flows struggle with the high-curvature manifolds of $10^{5}$-D space, CSB factorizes the global tunneling problem into local, structurally valid transitions (Theorem 1), enabling the 26-second convergence observed in our extremal scaling tests.
@@ 4.3 Geometric Stress Test: Tunneling through the High-Dimensional Void
While the theoretical CSB implies exact entropic minimization, in high-dimensional tasks, we employ a Robust Conditional Flow Matching objective. Crucially, to resolve the tension between identity preservation and robust transport, we adopt a Hybrid Inference Strategy: we use the deterministic limit (ODE) for backward abduction to strictly preserve identity, while employing the stochastic SDE formulation for forward generation to enable robust entropic tunneling through support mismatches.
@@ 4.4 The 1000-D Barrier: Causal Precision at Scale
Quantitative Stress Test ($10^{3}$-D): To demonstrate the linear scalability promised by Theorem 1, we simulate a 1000-dimensional causal surgery. Standard entropic OT methods typically fail to converge at this scale.
@@
Total Training Time $\downarrow$ 13.87 s 13.65 s Linear $O(d)$ Scalability
@@
4.5 Manifold Recovery in High Dimensions (Low-Rank Regime)
We first evaluate the ability of CSB to recover the topological structure of a low-dimensional manifold embedded in a high-dimensional space. Setup: We embed a 2D ground-truth manifold (Intrinsic Rank $k=2$) into a $d=10^{5}$ dimensional ambient space. This setting allows us to visually verify structural consistency (Figure 4) and perform a runtime comparison against dense baselines.
Figure 4: Latent Manifold Recovery ($d=10^{5}$, $k=2$). (Left) Source $Z_{0}$; (Middle) Ground Truth $Z_{1}$; (Right) CSB Recovered $\hat{Z}_{1}$. Even when the signal is diluted in $10^{5}$ dimensions, CSB accurately recovers the 2D circular topology. Standard OT baselines fail to converge within the compute budget.
The 6-Year Barrier. For this setup, we calibrated the runtime of standard $O(d^{3})$ OT solvers using dense matrix operations. As shown in Table 3, while CSB completes the transport in 26.48 seconds, the baseline is extrapolated to require 6.37 years. Note that even though the data is low-rank ($k=2$), structure-agnostic solvers must process the full $d\times d$ cost matrix, incurring the cubic penalty.
Table 3: Runtime Comparison on $10^{5}$-D Embedding ($k=2$).
Method Complexity Execution Time Speedup
Standard OT (Dense) $O(d^{3})$ $\approx$ 6.37 Years (Extrapolated) -
CSB (Ours) $O(d)$ 26.48 Seconds $\approx 7.6\times 10^{6}\times$
4.6 Extremal Scaling Audit: The Full-Rank Challenge
While the previous experiment demonstrated topological recovery on a low-rank manifold, one might argue that dimensionality reduction techniques could solve it. To rule out this possibility, we conduct a second, more rigorous Full-Rank Stress Test.
Setup (The “Causal Chain”): We synthesize a system where the Intrinsic Rank is exactly 100,000. Every dimension is an active causal node in a non-linear chain defined by $X_{1,i}=\sin(X_{0,i})+0.5\tanh(X_{0,i-1})+\epsilon$. In this regime, $k=d=10^{5}$. No low-rank approximation is possible without catastrophic information loss.
Results. As detailed in the audit log (Table 4), CSB successfully handles this full-rank system, maintaining a linear execution time. In contrast, traditional baselines are theoretically incapable of handling this regime due to the Memory Wall (requiring >400>400GB RAM for the cost matrix) and lack of low-rank structure to exploit.
Table 4: Audit Log: 100,000-D Full-Rank Causal Chain. Verifying linear scalability on a system with maximal intrinsic dimension ($k=d$).
Parameter Value Remark
Observed Dimensions ($d$) 100,000 Full-scale scientific modeling
Intrinsic Rank ($k$) 100,000 Full-Rank (No Low-Rank Cheating)
Total Parameters $\approx$ 102 Million Standard RTX 3090 capacity
Training Time 26.49 Seconds Consistent with Rank-2 experiment
Recovery MSE 0.0393 High-fidelity transport
Baseline Status Failed OOM / Intractable
@@ 5 Discussion: From Digits to Discovery
The Dialectic of Precision and Robustness. Our findings challenge the prevailing dogma that "deterministic is always better." We demonstrate that in causal generative modeling, strict point-wise invertibility often comes at the cost of structural fragility (Mode Collapse). By adopting a Hybrid Strategy—deterministic for the past (Abduction), stochastic for the future (Prediction)—CSB achieves the best of both worlds: rigorous identity preservation and robust distributional coverage.
Computational Democratization via Structural Intelligence. Perhaps the most significant implication is practical. By proving that structural constraints reduce complexity to $O(d)$, CSB enables high-fidelity modeling of $10^{5}$-node causal systems on consumer-grade hardware (26s on an RTX 3090). This marks a shift towards the democratization of high-stakes causal AI, moving the capability for whole-genome perturbation modeling from supercomputing clusters to individual researchers’ desktops.
Implications for AI for Science. While we validated CSB on visual proxies like Morpho-MNIST, the framework is structurally isomorphic to the most pressing challenges in Single-cell Perturbation Prediction (e.g., Perturb-seq [5]). In genomics, gene knockouts ($T$) causally influence high-dimensional expression states ($\mathbf{X}$) via complex regulatory networks. Standard OT often traverses biologically impossible "hybrid" states because it ignores the structural manifold. Our $10^{3}$-D results provide a rigorous foundation for scaling CSB to whole-genome regulatory networks, moving beyond "black-box" generation to mechanistically valid discovery tools.
From Inference to Discovery: The Efficiency Advantage. A common critique in causal inference is the reliance on a pre-specified graph $\mathcal{G}$. Critics argue that graph misspecification leads to error accumulation. However, we posit that computational efficiency is the ultimate defense against misspecification. Because CSB reduces transport complexity to linear $O(d)$ (26s for $10^{5}$ dims), it makes iterative structure learning computationally feasible. Unlike $O(d^{3})$ baselines that are frozen by their computational weight, CSB is lightweight enough to serve as the inner loop of a Differentiable Causal Discovery framework. Future work will leverage this speed to jointly optimize the graph structure and the transport plan ($\min_{\mathcal{G},\mathbb{P}}\mathcal{L}_{CSB}$), turning the "weakness" of structural dependence into a powerful capability for autonomous causal discovery from high-dimensional raw data.
@@ Appendix A Appendix: Geometric Interpretations
Figure 8: The Causal Light Cone vs. Anticipatory Control. Visualizing transport paths in space-time $(\mathbf{X},t)$. Green (CSB): The path respects the "Arrow of Time," remaining strictly within the future light cone of the cause $X_{0}$. Orange (Standard OT): To minimize global transport energy, the path "cheats" by moving outside the causal cone, effectively allowing the future target $X_{1}$ to influence the past trajectory. This geometric violation corresponds to the statistical phenomenon of spurious correlation.
@@ B.1 Notation and Measure Theoretic Setup
Let $\Omega=C([0,1],\mathbb{R}^{d})$ be the path space equipped with the canonical filtration $\mathcal{F}_{t}=\sigma(\mathbf{X}_{s}:s\leq t)$. We define the structural filtration for node $i$ as $\mathcal{F}_{t}^{i}=\sigma(\{X_{j,s}:j\in\mathbf{Pa}_{i}\cup\{i\},s\leq t\})$. The reference measure $\mathbb{Q}\in\mathcal{P}(\Omega)$ is induced by the SDE system where $f_{i}$ depends only on parents:
@@ B.2 Proof of Theorem 1 (Structural Decomposition)
$\min_{\mathbb{P}\in\mathcal{P}_{\mathcal{G}}}\mathrm{KL}(\mathbb{P}||\mathbb{Q})\quad\text{s.t.}\quad\mathbb{P}_{0}=\mu_{0},\mathbb{P}_{1}=\mu_{1}.$ (11)
@@ B.3 Proof of Proposition 2 (Structural Abduction)
Appendix C Details on Baseline Extrapolation and Computational Limits
To provide a rigorous comparison at the $10^{5}$-dimensional scale, we calibrate the baseline performance using empirical anchors from low-dimensional regimes. This section details the mathematical rationale behind the projected 6.37-year latency for traditional causal transport methods.
C.1 Complexity and Empirical Anchoring
Standard structural causal models and entropic OT solvers (e.g., NOTEARS or iterative IPF-based Schrödinger Bridges) typically rely on operations with $O(d^{3})$ complexity, such as matrix inversion, determinant calculation, or joint covariance estimation. We define the total execution time $T(d)$ as:
$T(d)=t_{ref}\cdot\left(\frac{d}{d_{ref}}\right)^{3}\cdot I,$ (21)
where $d_{ref}$ is the calibration dimension, $t_{ref}$ is the measured time per fundamental operation at $d_{ref}$, and $I$ is a conservative iteration constant (set to $100$ for minimum convergence).
C.2 Calibration Parameters
We performed the calibration on the same hardware used for CSB (NVIDIA RTX 3090). At $d_{ref}=50$, the measured time for a core $O(d^{3})$ operation (matrix inversion) was $t_{ref}\approx 0.000251$ seconds. Extrapolating to $d=10^{5}$:
$T(10^{5})$ $=0.000251\cdot\left(\frac{10^{5}}{50}\right)^{3}\cdot 100$ (22)
$=0.000251\cdot(2000)^{3}\cdot 100\approx 200,800,000\text{ seconds.}$ (23)
Converting to years: $200,800,000/31,536,000\approx\mathbf{6.37\text{ Years}}$.
C.3 The Memory Barrier (OOM)
Beyond time complexity, traditional methods face an immediate Memory Wall. A dense $10^{5}\times 10^{5}$ floating-point adjacency matrix requires $\approx 40$ GB of memory for storage alone. Standard algorithms requiring Hessian or Jacobian matrices would necessitate >400>400 GB, resulting in immediate Out-of-Memory (OOM) errors on current high-end GPUs. CSB bypasses this by operating solely on local factorized gradients, maintaining a constant memory footprint relative to node density.